OpenStudy (anonymous):

How can you graph y=x+7

OpenStudy (fibonaccichick666):

So, have you picked any x values yet?

OpenStudy (anonymous):

You need and x

TheSmartOne (thesmartone):

Or you could use Desmos Graphing Calculator...

OpenStudy (anonymous):

The homework im doing is Solve a system of equations by graphing and i have to graph y=x+7 how can i do it?

OpenStudy (fibonaccichick666):

well, have you ever graphed points before?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

would the cordinates be 1-7?

OpenStudy (anonymous):

(1,7) sorry

OpenStudy (fibonaccichick666):

so, if you put 0,1,-1 in for x, what y values will you get?

OpenStudy (confluxepic):

Go to this website and type in the equation. https://www.desmos.com/calculator

OpenStudy (anonymous):

I use that site all the time

OpenStudy (fibonaccichick666):

that's not solving the problem people!

OpenStudy (anonymous):

Anyways, it is in y=mx + b form (slope-intercept form).

OpenStudy (anonymous):

Set your y intercept on a graph to positive 7 (b).

OpenStudy (anonymous):

x is the slope. the slope is 1. so, just mark a rise and a run of 1, and draw a line through that.

OpenStudy (anonymous):

It is a positive slope, so the line will go upwards. it is positive because x is positive, not negative

OpenStudy (anonymous):

First. Understand that this makes a line. When you have an equation like so " y=x+7 " this makes a line on the y axis. But the point of intersection depends on the number on the other side of the line. http://www.wolframalpha.com/input/?i=y%3Dx%2B7 With this, find out the slope which is +7. The positive helps you decide whether it is ascending or descending. 

OpenStudy (anonymous):

Yep. thar ya go

OpenStudy (anonymous):

But the slope is +1. The y-int is +7.

OpenStudy (anonymous):

Just like with the example equation " y=m+x ", m is your point of intersection. A negative would have your line going down in numbers like so.... The number paired with x is your slope. please excuse the typo

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

No worries! Sorry for the confusion